What is X?
log base8 (2X-1) = 1/3

Answer by rajagopalan(174)

log base8 (2X-1) = 1/3
8 power 1/3=2x-1
but 8=2 power 3
2 power (3*1/3)=2x-1
2 power 1=2x-1
2=2x-1
2+1=2x
3=2x
x=3/2=1.5 or 3/2